In a survey, 30 people were asked how much they spent on their child's last birthday gift. The results were roughly bell-shaped with a mean of $36 and standard deviation of $6. Find the margin of error at a 80% confidence level.


Give your answer to two decimal places.


1
Expert's answer
2021-10-21T10:22:31-0400

Following is the formula for margin of error (MOE):

"MOE= Z\times\frac{\sigma}{\sqrt{n}}"

Where;

Z (critical value at 80% confidence level) = 1.28

σ (standard deviation) = $6

n (sample size) = 30

"MOE= 1.28\times\frac{6}{\sqrt{30}}"

MOE = 1.40